OpenStudy (anonymous):

Which equation represents the line passing through the points (3, 2) and (−9, 6)? a. x − 3y = 9 b. x + 3y = 9 c. 3x − y = −9 d. 3x + y = 9

OpenStudy (anonymous):

@iGreen @geekfromthefutur

OpenStudy (anonymous):

please help...

OpenStudy (geekfromthefutur):

Ik what it is

OpenStudy (johnweldon1993):

First we need to find the slope using the formula \(\large m = \frac{y_2 - y_1}{x_2 - x_1}\) So we have \[\large m = \frac{6 - 2}{-9 - 3} = \frac{4}{-12} = -\frac{1}{3}\] So our slope is \(\large -\frac{1}{3}\) Now that we have a slope we can plug it (and 1 of our points up there) into the point slope formula \[\large y - y_0 = m(x - x_0)\] So this would look like (using the point (3,2)) \[\large y - 2 = -\frac{1}{3}(x - 3)\] Lets simplify it \[\large y = -\frac{1}{3}x + 3\] Now it may look like none of those lines are correct, but if we multiply this entire equation by 3...we have \[\large 3y = -x + 9\] Now lets just add 'x' to both sides and we will have \[\large 3y + x = 9\]

OpenStudy (anonymous):

thanks!!!

OpenStudy (johnweldon1993):

Anytime!